# Explainable Differential Privacy-Hyperdimensional Computing for Balancing Privacy and Transparency in Additive Manufacturing Monitoring

### v3

@@ 4.2 High-Speed Camera Data Collection
In the pursuit of advancing in-situ monitoring within the LPBF process, our experimental framework integrates a high-speed visible spectrum video camera, characterized by its rapid frame rate and high-resolution capabilities. Positioned strategically within the build chamber, the camera captures detailed thermal patterns as the laser fuses powder layers, creating a dynamic record of the melting process. The specificity of the camera’s placement and settings allows for an unprecedented view into the thermal dynamics at play, particularly in the formation of overhang features and the solidification of bulk areas.
@@ 4.3 Thermal Signature Classification
Building on the data captured, the study introduces a nuanced framework that delineates eight classes based on specific overhang and bulk characteristics across four distinct strips. This classification is not merely academic; it has profound implications for understanding the thermal behavior in AM and the structural integrity and surface quality of the printed parts. Each class represents a unique thermal signature, corresponding to different combinations of overhang features and bulk regions, thereby enabling a detailed analysis of potential defects and irregularities.
@@ 4.4 Privacy-Enhanced Additive Manufacturing
As we navigate through the complexities of in-situ sensing and data analysis, the imperative of preserving privacy emerges as a critical concern. The integration of DP techniques in the processing of sensor data addresses this concern head-on, ensuring that while the data provides invaluable insights into the LPBF process, it does not expose sensitive information related to part designs or proprietary manufacturing techniques. This dual focus on enhancing manufacturing precision while safeguarding data privacy forms the cornerstone of our approach, setting a new standard for research and practice in the field of AM.
@@ 5 Experimental Results
In this section, we thoroughly evaluate the DP-HD framework, focusing on both training and inference privacy. We begin by exploring encoding strategies, and examining the impact of exclusive and inclusive methods in mapping optical images to hyperspace. We then analyze how adding noise to class hypervectors affects the model’s memorization capabilities and determine the optimal SD for generating random basis vectors. Additionally, we assess the explainability of the DP-HD model and benchmark its performance against the state-of-the-art DP-SGD mechanism using Opacus, an open-source library for PyTorch that integrates DP into deep learning workflows [79]. For inference privacy, we demonstrate the decoding of hypervectors to infer optical images from the LPBF process and present results on safeguarding query hypervectors. Finally, we compare the inference time of our proposed model with leading ML models in image classification tasks, highlighting the importance of efficient inference in in-situ sensing systems.
@@ 5.1.1 Sensitivity and Uncertainty Analysis
5.1.2 Benchmark Comparison
In this analysis, we benchmarked our newly developed DP-HD model against established models in the field of image classification. This comparison includes an evaluation alongside models such as ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. To implement DP, we utilized Opacus, a freely available open-source library designed for PyTorch that facilitates training deep learning models with DP mechanisms. Table 4 details the accuracies post-DP application, revealing that DP-HD sustains a minor accuracy reduction compared to significant drops in other models. With a stringent privacy level, represented by an $\epsilon$ value of 0.6, DP-HD maintains an accuracy of 91.04%, whereas other models plummet below 43%, showcasing DP-HD’s robustness against high noise levels as it distributes memory evenly across all dimensions of the hyperspace. As $\epsilon$ increases to 0.8, reducing the noise, DP-HD’s accuracy climbs to 93.34%, significantly surpassing other models which remain under 54%. This performance gap widens further when $\epsilon$ is adjusted to 1, with DP-HD reaching an accuracy of 94.43%, and the closest competitor, DenseNet201, achieves only 69.13%, still 25.3% lower than DP-HD. As $\epsilon$ increases further to 2, the accuracy of DP-HD is 96.30%, while the other models’ accuracies—75.22% for ResNet50, 82.48% for GoogLeNet, 58.20% for AlexNet, 77.70% for DenseNet201, and 57.20% for EfficientNet B2—illustrate a substantial disparity, with GoogLeNet being the closest yet still 13.82% less accurate than DP-HD. Even at an $\epsilon$ of 3, where privacy levels are lower and noise reduction is significant, DP-HD’s accuracy remains high at 96.66%, outperforming ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2, which exhibit accuracies of 81.18%, 83.73%, 56.27%, 86.89%, and 56.46% respectively. This demonstrates DP-HD’s superior capability to balance privacy and performance compared to other prevalent ML models in the field.
@@ 5.2 Inference Privacy
Figure 13: Inference time for defect detection of 1610 query samples.
@@ 6 Discussion
The choice of exclusive versus inclusive encoding in HD also impacts the balance between accuracy and privacy. As our results indicate, encoding choices should adapt based on the distribution within each class and the distance between class centers to maintain high accuracy and privacy. Importantly, our findings reveal that the optimal standard deviation of random basis vectors remains stable across different HD configurations, as long as the dataset distribution is adequately represented. This stability further reinforces that our framework’s reliability does not depend on specific HD model parameters or sample sizes, enhancing its robustness.
@@ 7 Conclusion and Future Works
This paper introduced the Differential Privacy-Hyperdimensional Computing (DP-HD) framework as an innovative solution for addressing privacy concerns in Additive Manufacturing (AM). By integrating Differential Privacy (DP) mechanisms within the Hyperdimensional Computing (HD) model, the DP-HD framework ensures the protection of sensitive data during training and inference while maintaining high model performance. The key advantage of DP-HD is its explainability, enabling the prediction of noise impact on model accuracy, which is critical for in-situ sensing applications in AM. Our experiments demonstrated that DP-HD preserves privacy without significantly compromising accuracy, showcasing its robustness and suitability for real-time monitoring in AM. The ability to balance privacy and performance through explainable insights sets DP-HD apart, making it a valuable tool for enhancing manufacturing precision while safeguarding data privacy. While the DP-HD framework demonstrates significant promise, certain aspects warrant further exploration to enhance its applicability. Currently, finding the optimal encoding strategy relies on selecting training and validation subsets. Future work focuses on developing mathematical methods to automatically determine the best Standard Deviation (SD) for generating random basis vectors based on data distribution. Additionally, due to DP-HD’s success in privacy-preserving techniques, the framework can be integrated within a federated learning structure to leverage distributed data processing while maintaining robust privacy measures. Future work can explore applying the DP-HD framework in other sensitive domains, such as healthcare, financial services, and government data management, where balancing privacy and accuracy is essential for secure and effective machine learning applications.
@@ Acknowledgment
This work was supported by the National Science Foundation [grant numbers 2127780, 2312517]; the Semiconductor Research Corporation (SRC); the Office of Naval Research [grant numbers N00014-21-1-2225, N00014-22-1-2067]; the Air Force Office of Scientific Research [grant number FA9550-22-1-0253]; UConn Startup Funding, and generous gifts from Xilinx and Cisco. The authors gratefully acknowledge the valuable contributions of the National Institute of Standards and Technology (NIST), particularly Dr. Brandon Lane, who provided data for this research.

### v4

@@ 4.2 High-Speed Camera Data Collection
In the pursuit of advancing in-situ monitoring within the LPBF process, our experimental framework integrates a high-speed visible spectrum video camera, characterized by its rapid frame rate and high-resolution capabilities. Positioned strategically within the build chamber, the camera captures detailed thermal patterns as the laser fuses powder layers, creating a dynamic record of the melting process. This setup provides a realistic basis for investigating defect formation mechanisms and validating the effectiveness of DP-HD in industrially relevant scenarios. The specificity of the camera’s placement and settings allows for an unprecedented view into the thermal dynamics at play, particularly in the formation of overhang features and the solidification of bulk areas.
@@ 4.3 Thermal Signature Classification
Building on the data captured, the study introduces a nuanced framework that delineates eight classes based on specific overhang and bulk characteristics across four distinct strips. This classification is not merely academic; it provides actionable insights into industrial AM challenges, such as identifying defects that could compromise part quality or structural integrity. Each class represents a unique thermal signature, corresponding to different combinations of overhang features and bulk regions, thereby enabling a detailed analysis of potential defects and irregularities.
@@ 4.4 Privacy-Enhanced Additive Manufacturing
As we navigate through the complexities of in-situ sensing and data analysis, the imperative of preserving privacy emerges as a critical concern. The integration of DP techniques in the processing of sensor data addresses this concern head-on, ensuring that while the data provides invaluable insights into the LPBF process, it does not expose sensitive information related to part designs or proprietary manufacturing techniques. This integration is particularly relevant for industries such as aerospace and automotive manufacturing, where data privacy is paramount. This dual focus on enhancing manufacturing precision while safeguarding data privacy forms the cornerstone of our approach, setting a new standard for research and practice in the field of AM.
@@ 5 Experimental Results
In this section, we thoroughly evaluate the DP-HD framework, focusing on both training and inference privacy. We begin by exploring encoding strategies, and examining the impact of exclusive and inclusive methods in mapping optical images to hyperspace. We then analyze how adding noise to class hypervectors affects the model’s memorization capabilities and determine the optimal SD for generating random basis vectors. Additionally, we assess the explainability of the DP-HD model and benchmark its performance against the state-of-the-art DP-SGD and RDP mechanisms, implemented using Opacus, an open-source library for PyTorch that integrates DP into deep learning workflows [83]. For inference privacy, we demonstrate the decoding of hypervectors to infer optical images from the LPBF process and present results on safeguarding query hypervectors. Finally, we compare the inference time of our proposed model with leading ML models in image classification tasks, highlighting the importance of efficient inference in in-situ sensing systems.
@@ 5.1.1 Sensitivity and Uncertainty Analysis
5.1.2 Benchmark Comparison
In this study, we benchmarked our newly developed DP-HD model against established models in the field of image classification. This comparison includes an evaluation alongside models such as ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2. To implement DP-SGD, we utilized Opacus, a freely available open-source library designed for PyTorch that facilitates training deep learning models with DP mechanisms. Table 4 details the accuracies post-DP application, revealing that DP-HD sustains a slight accuracy reduction compared to significant drops in other models. With a stringent privacy level, represented by an $\epsilon$ value of 0.6, DP-HD maintains an accuracy of 91.04%, whereas other models, securing using DP-SGD, plummet below 43%, showcasing DP-HD’s robustness against high noise levels as it distributes memory evenly across all dimensions of the hyperspace. As $\epsilon$ increases to 0.8, reducing the noise, DP-HD’s accuracy climbs to 93.34%, significantly surpassing other models which remain under 54%. This performance gap expands further when $\epsilon$ is adjusted to 1, with DP-HD reaching an accuracy of 94.43%, and the closest competitor, DenseNet201, achieves only 69.13%, still 25.3% lower than DP-HD. As $\epsilon$ increases further to 2, the accuracy of DP-HD is 96.30%, while the other models’ accuracies—75.22% for ResNet50, 82.48% for GoogLeNet, 58.20% for AlexNet, 77.70% for DenseNet201, and 57.20% for EfficientNet B2—illustrate a substantial disparity, with GoogLeNet being the closest yet still 13.82% less accurate than DP-HD. Even at an $\epsilon$ of 3, where privacy levels are lower and noise reduction is significant, DP-HD’s accuracy remains high at 96.66%, outperforming ResNet50, GoogLeNet, AlexNet, DenseNet201, and EfficientNet B2, which exhibit accuracies of 81.18%, 83.73%, 56.27%, 86.89%, and 56.46% respectively. This demonstrates DP-HD’s superior capability to balance privacy and performance compared to other popular ML models in the field.
@@ 5.1.2 Benchmark Comparison
RDP ResNet50 31.55% 36.71% 61.37% 75.78% 88.94%
RDP GoogLeNet 12.48% 38.45% 40.99% 73.42% 79.69%
RDP AlexNet 34.66% 57.70% 52.92% 55.03% 55.59%
RDP DenseNet201 29.19% 17.02% 68.26% 66.15% 83.17%
RDP EfficientNet B2 29.69% 44.66% 47.08% 61.24% 63.98%
For the comparison of DP-HD with RDP, Table 4 clearly demonstrates the superior performance of DP-HD across all privacy budget values. At the lowest privacy budget ($\epsilon=0.6$), DP-HD achieves an accuracy of 91.04%, while the highest-performing RDP-based model, AlexNet, reaches only 34.66%, resulting in a gap of 56.38%. This substantial difference highlights DP-HD’s robustness in maintaining high accuracy even under strict privacy constraints. As $\epsilon$ increases to 0.8, DP-HD’s accuracy improves to 93.34%, whereas AlexNet, the best RDP-based model, only reaches 57.70%, maintaining a difference of 35.64%. At $\epsilon=1$, DP-HD achieves 94.43%, significantly outperforming DenseNet201 (68.26%) and other RDP-based models, such as ResNet50 (61.37%) and AlexNet (52.92%), with at least a 26.17% margin over the best competitor. At higher privacy budgets, such as $\epsilon=2$, DP-HD maintains its dominance with an accuracy of 96.30%, compared to ResNet50’s 75.78%, DenseNet201’s 66.15%, and GoogLeNet’s 73.42%. The gap between DP-HD and the top-performing RDP-based model remains significant, at 20.52%. Finally, at $\epsilon=3$, DP-HD achieves a remarkable accuracy of 96.66%, surpassing ResNet50 (88.94%) and DenseNet201 (83.17%) by 7.72% and 13.49%, respectively. This consistent outperformance across all privacy budgets emphasizes DP-HD’s ability to balance high accuracy and strong privacy guarantees more effectively than RDP-based implementations.
When evaluating DP-HD against both DP-SGD and RDP methods across varying training set sizes, with the privacy budget fixed at $\epsilon=1$, DP-HD continues to showcase superior performance. As seen in Figure 11(a), and 11(b), DP-HD achieves an accuracy of 73.26% at the smallest training size of 1,000 samples, significantly outperforming ResNet50 with DP-SGD (13.66%) and RDP (14.00%). This trend persists as the training size increases, with DP-HD maintaining its lead across all benchmarks. At a training size of 3,500 samples, DP-HD reaches 94.02% accuracy, while the closest competitor, AlexNet using DP-SGD, achieves 61.37%. RDP-based AlexNet follows with 62.88%, still trailing by over 31.14%.
The robustness of DP-HD becomes even more apparent when comparing its consistency across different training set sizes. Models using DP-SGD, such as ResNet50 and GoogLeNet, exhibit fluctuating accuracies, often struggling to surpass 40% at smaller training sizes. Similarly, RDP-based models, while slightly more stable, still fail to match the consistent improvements of DP-HD as the training size grows. These results highlight DP-HD’s ability to leverage its hyperdimensional representation for enhanced accuracy and stability, solidifying its position as a robust solution for privacy-preserving ML tasks.
Figure 11: Defect detection accuracy across varying training sample sizes, comparing the proposed DP-HD method with (a) DP-SGD and (b) RDP-based benchmarks.
@@ 6 Discussion
The choice of exclusive versus inclusive encoding in HD also impacts the balance between accuracy and privacy. As our results indicate, encoding choices should adapt based on the distribution within each class and the distance between class centers to maintain high accuracy and privacy. Importantly, our findings reveal that the optimal SD of random basis vectors remains stable across different HD configurations, as long as the dataset distribution is adequately represented. This stability further reinforces that our framework’s reliability does not depend on specific HD model parameters or sample sizes, enhancing its robustness.
Additionally, the only hyperparameter requiring manual selection in the DP-HD framework is the size of the hypervectors, which must be sufficiently large to capture all essential information from the feature vectors. For our experimental setup, we selected a hypervector size of 1000, as this was the minimum size required to preserve relevant information effectively. To validate this choice, we tested the framework with higher hypervector sizes and found that while larger sizes resulted in marginal accuracy improvements, the gains were negligible and did not justify the additional computational cost. This demonstrates the robustness of HD-based models, which maintain stable performance as long as the hypervector size exceeds a certain threshold. In the case of DP-HD, this threshold was determined to be 1000, ensuring reliable performance without extensive hyperparameter tuning.
@@
For future work, the proposed DP-HD framework can be extended to other domains that require a balance between privacy and accuracy in ML models trained on sensitive data. While this study focuses on AM, its utility is particularly relevant for privacy-sensitive industrial environments such as aerospace, automotive, and energy production. For instance, real-time defect detection in these sectors often involves proprietary designs and sensitive operational data, where privacy breaches could have financial and regulatory consequences. The DP-HD framework provides a scalable solution to address these challenges while maintaining high accuracy. DP-HD can also be applied in sectors such as healthcare, financial services, and government data management. These domains frequently involve training ML models on confidential information, where ensuring data privacy is crucial while maintaining high model performance.
In healthcare, ML models are often trained on sensitive patient data, including medical histories and diagnostic records, which are vulnerable to privacy breaches. Financial services use ML models for tasks like fraud detection and risk assessment, which rely on sensitive customer information, such as transaction records and credit history. Government data management can involve ML models analyzing citizen data, including personal identification and behavioral records, which are susceptible to misuse if privacy is compromised. Applying the DP-HD framework in these areas can help protect confidential information while ensuring that model accuracy remains effective for these critical applications.
Figure 14: Inference time for defect detection of 1610 query samples.
@@ 7 Conclusion and Future Works
This paper introduced the Differential Privacy-Hyperdimensional Computing (DP-HD) framework as an innovative solution for addressing privacy concerns in Additive Manufacturing (AM). By integrating Differential Privacy (DP) mechanisms within the Hyperdimensional Computing (HD) model, the DP-HD framework ensures the protection of sensitive data during training and inference while maintaining high model performance. The key advantage of DP-HD is its explainability, enabling the prediction of noise impact on model accuracy, which is critical for in-situ sensing applications in AM. Our experiments demonstrated that DP-HD preserves privacy without significantly compromising accuracy, showcasing its robustness and suitability for real-time monitoring in AM. The ability to balance privacy and performance through explainable insights sets DP-HD apart, making it a valuable tool to improve manufacturing precision while safeguarding data privacy. While the DP-HD framework demonstrates significant promise, certain aspects warrant further exploration to enhance its applicability. Currently, finding the optimal encoding strategy is based on the selection of training and validation subsets. Future work focuses on developing mathematical methods to automatically determine the best Standard Deviation (SD) for generating random basis vectors based on data distribution. Moreover, deploying DP-HD in collaboration with industry partners can validate its scalability and adaptability for practical industrial applications, such as real time defect detection in aerospace and automotive manufacturing, where privacy concerns often intersect with stringent performance requirements. Additionally, due to DPHD’s success in privacy preserving techniques, the framework can be integrated within a federated learning structure to leverage distributed data processing while maintaining robust privacy measures. Beyond AM, the DP-HD framework holds considerable potential in other domains requiring secure and efficient machine learning applications. Future work will explore its applicability in areas such as healthcare, financial services, and government data management. In these fields, the ability to maintain a delicate balance between data privacy and model performance is crucial for safeguarding sensitive information while delivering reliable, high quality predictions.
@@ Acknowledgment
This work was supported by the National Science Foundation, United States [grant numbers 2127780, 2312517, 2434519]; the Semiconductor Research Corporation (SRC), United States; the Office of Naval Research, United States [grant numbers N00014-21-1-2225, N00014-22-1-2067]; the Air Force Office of Scientific Research, United States [grant number FA9550-22-1-0253]; UConn Startup Funding, and generous gifts from Xilinx and Cisco. The authors gratefully acknowledge the valuable contributions of the National Institute of Standards and Technology (NIST, United States), particularly Dr. Brandon Lane, who provided data for this research.